Calculate the pH of the following solution:
0.1 M HF + 0.2 M NaCN
Given: pKa for: HF/F- = 3.2 HCN/CN-= 9.4
1
Expert's answer
2013-03-29T12:26:53-0400
The chemical equation for this reaction is

HF + NaCN = NaF + HCN

HCN is much more weak acid than HF because it's pKa is greater

If the volumes of HF and NaCN were equal, HF is limiting reagent and is completely transformed into NaF. In the solution there is buffer HCN/CN-, and the pH of this buffer is

pH = pKa + lg[A-]/[HA]
pH = 9.4 + lg([CN-]/([HCN]) = 9.4 + lg (0.1/0.1) = 9.4
Answer: pH = 9.4